# Deepfake Generation and Detection: A Benchmark and Survey

### v3

@@ 3.1.2 Face Reenactment
DR2 [19] WACV’24 Lack emotional and other latent attributes control. The model explored effective strategies for reducing the training workload.
@@
Multimodal PC-AVS [244] CVPR’21 ➊➋ Lack emotional and other latent attributes control. Introduction of pose-source video drive compensation to generate head motion in video.
@@
GC-AVT [245] CVPR’22 ➋➌ Poor resolution, unable to handl complex backgrounds. In addition to the source image, a gesture source, an expression source, and audio are introduced to drive the talking head generation.
@@
Yu et al[246] TMM’22 Lack emotional and other latent attributes control. Fusion of audio and text inputs for more accurate lip movement and chin posture prediction.
@@
Xu et al[247] CVPR’23 Insufficient control over the intensity of emotional output. Embedding textual, visual, and auditory emotional modalities into a unified space.
@@
LipFormer [248] CVPR’23 ➎➓ Poor ability to preserve face feature attributes. Propose retaining high-quality facial details obtained from pre-training in a codebook format and repro -ducing them by driving the encoded mapping relationship between audio and lip movements.
@@
Wang et al[249] TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
@@
Diffusion DAE-Talker [250] MM’23 High model complexity. It replaces traditional manually crafted intermediate representations with data-driven latent representa -tions obtained from a DAE.
@@
Yu et al[251] ICCV’23 ➋➒ Poor resolution, high model complexity. Building a corresponding mapping between audio and non-lip representations and training using the diffusion model.
@@
Stypułkowski [252] WACV’24 ➊➑ High model complexity, short video generation duration. The model incorporates motion frame and audio embedding information to capture past movements and future expressions, with an emphasis on the mouth region through an additional lip sync loss.
@@
EmoTalker [253] ICASSP’24 ➌➑ High model complexity. It achieves emotion-editable talking face generation based on a conditional diffusion model.
@@
3D-Model AD-NeRF [254] ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
@@
DFRF [255] ECCV’22 Lack of emotional and other latent attributes control. Combining audio with 3D perceptual features and proposing an facial deformation module.
@@
AE-NeRF [256] AAAI’24 Lack emotional and other latent attributes control. Facial modeling is divided into NeRF related to audio and unrelated to audio to enhance audio-visual lip synchronization and facial detail.
@@
SyncTalk [257] CVPR’24 Lack controllable emotional intensity regulation. The facial sync controller boosts component coordination, and a portrait generator corrects artifacts, enhancing video details.
@@
Ye et al[258] ICLR’24 [76] Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.
@@ 3.1.3 Talking Face Generation
In this section, we review current methods from three perspectives: audio/text driven, multimodal conditioned, diffusion-based, and 3D-model Technologies. We also summarize them in Tab. IV.
@@
$\bullet$ Audio/Text Driven. Methods aim to map and guide lip and facial movements in generated videos by understanding the semantic information from the driving source [259, 255, 243]. Early methods [119, 105] perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers [120, 106] propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk [238] decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. SadTalker [241] extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods [260, 97, 261, 262] propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos [263, 240]. Recently, some methods [264, 265, 266] extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN [264] establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO [265] employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP [267] introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et al[268] propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 [19] explores practical strategies for reducing the training workload.
@@
$\bullet$ Multimodal Conditioned. To generate more realistic talking videos, some methods [244, 245, 247, 248] introduce additional modal information on top of audio-driven methods to guide facial pose and expression. GC-AVT [245] generates realistic talking videos by independently controlling head pose, audio information, and facial expressions. This approach introduces an expression source video, providing emotional information during the speech and the pose source video. However, the video quality falls below expectations, and it struggles to handle complex background changes. Xu et al[247] integrate text, image, and audio-emotional modalities into a unified space to complement emotional content in textual information. Multimodal approaches have significantly enhanced the vividness of generated videos, but there is still room for exploration of organically combining information driven by different sources and modalities.
@@
$\bullet$ Diffusion-based. Recently, some methods [17, 18, 250, 269, 253] apply the Diffusion model to the task of talking face generation. For fine-grained talking video generation, DAE-Talker [250] replaces manually crafted intermediate representations, such as facial landmarks and 3DMM coefficients, with data-driven latent representations obtained from a Diffusion Autoencoder (DAE). The image decoder generates video frames based on predicted latent variables. EmoTalker [253] utilizes a conditional diffusion model for emotion-editable talking face generation. It introduces emotion intensity blocks and the FED dataset to enhance the model’s understanding of complex emotions. Very recently, diffusion models are gaining prominence in talking face generation tasks [252] and video generation tasks [148, 270].
@@
$\bullet$ 3D-model Technologies. 3D models, exemplified by NeRF, are gaining traction in talking face generation [254, 255, 256]. AD-NeRF [254] directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic NeRF. AE-NeRF [256] employs a dual NeRF framework to separately model audio-related regions and audio-independent regions. Furthermore, some methods [257, 258] adopt Tri-Plane [271] to represent facial and audio attributes. Synctalk [257] models and renders head motion using a tri-plane hash representation, and then further enhances the output quality using a portrait synchronization generator. In the foreseeable future, 3D Gaussian Splatting [272] will also be widely applied to this task.
@@
TABLE V: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.). Full table can be viewed in Appendix B.
@@
GeneGAN [273] BMVC’17 G. ➋➓ In early attribute editing, separate models were trained for specific attributes. The key idea was to reassemble attribute vectors in the latent space, achieving successful editing.
@@
SC-FEGAN [274] ICCV’19 G. Users can generate high-quality edited output images by freely sketching parts of the source image.
@@
Shen et al[275] CVPR’20 G. Thoroughly investigated how to encode different semantics in the latent space and explored the disentanglement between var -ious semantics to achieve precise control over facial attributes.
@@
Yao et al[276] ICCV’21 G.+T. By integrating explicit decoupling terms and identity-consistent terms into the loss function, the preservation of facial identity information is improved, resulting in high-quality face editing in videos.
@@
HifaFace [277] CVPR’21 G. Proposed a solution based on wavelet transform to address the issue of partial loss of attribute information when generating edited results due to "cyclic consistency" problems.
@@
Preechakul et al[278] CVPR’22 D. When encoding images, it is divided into semantically meaningful parts and parts that represent the details of the image.
@@
FENeRF [279] CVPR’22 G.+NeRF ➊➍ The introduction of semantic masks into the conditional radiance field enables finer image textures.
@@
FDNeRF [29] SIG.’22 G.+NeRF The introduction of the Conditional Feature Warping (CFW) module addresses the issue of temporal inconsistency caused by dynamic information in the process of face editing in videos.
@@
AnyFace [20] CVPR’22 G. ➏➑ Proposed a dual-branch framework for text-driven facial editing, with coordination achieved between the two branches through a Cross-Modal Distillation (CMD) module.
@@
Huang et al[280] CVPR’23 D. Proposed the concept of assisted diffusion, integrating individual multimodalities to explore the complementarity between different modalities.
@@
Ozkan et al[281] ICCV’23 G. The entangled attribute space is decomposed into conceptual and hierarchical latent spaces, and transformer network encoders are employed to modify information in the latent space.
@@
CIPS-3D++ [282] TPAMI’23 G.+NeRF ➊➒ Replaced the convolutional architecture with an MLP (Multi-Layer Perceptron) architecture to achieve faster rendering speeds.
@@
TG-3DFace [283] ICCV’23 G. ➌➏ For different scenarios, two sets of text-to-face cross-modal alignment methods were designed with specific focuses.
@@
VecGAN++ [284] TPAMI’23 G. Orthogonal constraint and disentanglement loss are used to decouple attribute vectors in the latent space.
@@
DiffusionRig [285] CVPR’23 D. 3DMM and diffusion model integration propose a two-stage method for learning personalized facial details.
@@
Kim et al[286] CVPR’23 D. Proposed a method for facial editing in videos based on the diffusion model.
@@
SDGAN et al[287] AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
@@
FaceDNeRF [288] NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.

### v4

@@ 3.1.2 Face Reenactment
DR2 zhang2024dr2 WACV’24 Lack emotional and other latent attributes control. The model explored effective strategies for reducing the training workload.
RADIO lee2024radio WACV’24 ➊➋➏ Lack emotional and other latent attributes control. Introducing StyleGAN2 style modulation to adapt to human identity and utilizes ViT blocks to focus on facial attributes in the reference image.
@@
Multimodal PC-AVS zhou2021pose CVPR’21 ➊➋ Lack emotional and other latent attributes control. Introduction of pose-source video drive compensation to generate head motion in video.
@@
GC-AVT liang2022expressive CVPR’22 ➋➌ Poor resolution, unable to handl complex backgrounds. In addition to the source image, a gesture source, an expression source, and audio are introduced to drive the talking head generation.
@@
Yu et alyu2021multimodal TMM’22 Lack emotional and other latent attributes control. Fusion of audio and text inputs for more accurate lip movement and chin posture prediction.
@@
Xu et alxu2023high CVPR’23 Insufficient control over the intensity of emotional output. Embedding textual, visual, and auditory emotional modalities into a unified space.
@@
LipFormer wang2023lipformer CVPR’23 ➎➓ Poor ability to preserve face feature attributes. Propose retaining high-quality facial details obtained from pre-training in a codebook format and repro -ducing them by driving the encoded mapping relationship between audio and lip movements.
@@
Wang et alwang2024styletalk++ TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
@@
Diffusion DAE-Talker du2023dae MM’23 High model complexity. It replaces traditional manually crafted intermediate representations with data-driven latent representa -tions obtained from a DAE.
@@
Yu et alyu2023talking ICCV’23 ➋➒ Poor resolution, high model complexity. Building a corresponding mapping between audio and non-lip representations and training using the diffusion model.
DreamTalk ma2023dreamtalk arXiv’23 ➋➌➏ Mismatched emotion and semantics, occasional artifacts. The denoising network, style-aware lip expert, and style predictor collaborate to make the model perform well in various speaking styles.
@@
Stypułkowski stypulkowski2024diffused WACV’24 ➊➑ High model complexity, short video generation duration. The model incorporates motion frame and audio embedding information to capture past movements and future expressions, with an emphasis on the mouth region through an additional lip sync loss.
@@
EmoTalker zhang2024emotalker ICASSP’24 ➌➑ High model complexity. It achieves emotion-editable talking face generation based on a conditional diffusion model.
VASA-1 xu2024vasa arXiv’24 High model complexity. Expressive and well-decoupled facial latent space has been constructed, and highly controllable, high -quality generation effects have been achieved based on the Diffusion Transformer.
@@
3D-Model AD-NeRF guo2021ad ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
@@
DFRF shen2022learning ECCV’22 Lack of emotional and other latent attributes control. Combining audio with 3D perceptual features and proposing an facial deformation module.
@@
AE-NeRF li2023ae AAAI’24 Lack emotional and other latent attributes control. Facial modeling is divided into NeRF related to audio and unrelated to audio to enhance audio-visual lip synchronization and facial detail.
@@
SyncTalk peng2023synctalk CVPR’24 Lack controllable emotional intensity regulation. The facial sync controller boosts component coordination, and a portrait generator corrects artifacts, enhancing video details.
@@
Ye et alye2024real3d ICLR’24 zhu2022celebv Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.
@@ 3.1.3 Talking Face Generation
In this section, we review current methods from three perspectives: audio/text driven, multimodal conditioned, diffusion-based, and 3D-model Technologies. We also summarize them in Table 4.
@@
$\bullet$ Audio/Text Driven. Methods aim to map and guide lip and facial movements in generated videos by understanding the semantic information from the driving source apb2face ; shen2022learning ; wang2023seeing . Early methods fan2015photo ; chen2018lip perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers chen2019hierarchical ; prajwal2020lip propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk zhou2020makelttalk decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. SadTalker zhang2023sadtalker extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods wu2023speech2lip ; wang2023memory ; lee2024radio ; fu2023mimic propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos sheng2023stochastic ; gan2023efficient . Recently, some methods tan2023emmn ; zhai2023talking ; gururani2023space extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN tan2023emmn establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO zhai2023talking employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP ma2023talkclip introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et alzhong2023identity propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 zhang2024dr2 explores practical strategies for reducing the training workload.
@@
$\bullet$ Multimodal Conditioned. To generate more realistic talking videos, some methods zhou2021pose ; liang2022expressive ; xu2023high ; wang2023lipformer introduce additional modal information on top of audio-driven methods to guide facial pose and expression. GC-AVT liang2022expressive generates realistic talking videos by independently controlling head pose, audio information, and facial expressions. This approach introduces an expression source video, providing emotional information during the speech and the pose source video. However, the video quality falls below expectations, and it struggles to handle complex background changes. Xu et alxu2023high integrate text, image, and audio-emotional modalities into a unified space to complement emotional content in textual information. Multimodal approaches have significantly enhanced the vividness of generated videos, but there is still room for exploration of organically combining information driven by different sources and modalities.
@@
$\bullet$ Diffusion-based. Recently, some methods mir2023dit ; ma2023dreamtalk ; du2023dae ; xu2023multimodal ; zhang2024emotalker apply the Diffusion model to the task of talking face generation. For fine-grained talking video generation, DAE-Talker du2023dae replaces manually crafted intermediate representations, such as facial landmarks and 3DMM coefficients, with data-driven latent representations obtained from a Diffusion Autoencoder (DAE). The image decoder generates video frames based on predicted latent variables. EmoTalker zhang2024emotalker utilizes a conditional diffusion model for emotion-editable talking face generation. It introduces emotion intensity blocks and the FED dataset to enhance the model’s understanding of complex emotions. Very recently, diffusion models are gaining prominence in talking face generation tasks stypulkowski2024diffused ; wei2024aniportrait ; tian2024emo and video generation tasks wang2023leo ; karras2023dreampose . Emo tian2024emo directly predicts video from audio without the need for intermediate 3D components, achieving excellent results. However, the lack of explicit control signals may easily lead to unnecessary artifacts. Based on the Diffusion Transformer architecture, VASA-1 xu2024vasa finely encodes and reconstructs facial details, constructing an expressive and well-decoupled facial latent space. This results in highly controllable and high-resolution video generation.
@@
$\bullet$ 3D-model Technologies. 3D models, exemplified by NeRF, are gaining traction in talking face generation guo2021ad ; shen2022learning ; li2023ae . AD-NeRF guo2021ad directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic NeRF. AE-NeRF li2023ae employs a dual NeRF framework to separately model audio-related regions and audio-independent regions. Furthermore, some methods peng2023synctalk ; ye2024real3d adopt Tri-Plane chan2022efficient to represent facial and audio attributes. Synctalk peng2023synctalk models and renders head motion using a tri-plane hash representation, and then further enhances the output quality using a portrait synchronization generator. Very recently, 3D Gaussian Splatting kerbl20233d also been widely applied to this task. Some method chen2024gstalker ; cho2024gaussiantalker ; li2024talkinggaussian ; yu2024gaussiantalker introduce 3DGS to achieve more refined facial reconstruction and motion details, aiming to address the issue of insufficient pose and expression control caused by NeRF’s implicit representation.
@@
Table 5: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.).
@@
SC-FEGAN jo2019sc ICCV’19 G. Users can generate high-quality edited output images by freely sketching parts of the source image.
AttGAN he2019attgan TIP’19 G. Applying attribute classification constraints to generated images has validated the drawbacks of enforcing stringent attribute independence constraints in latent representations.
@@
Shen et alshen2020interpreting CVPR’20 G. Thoroughly investigated how to encode different semantics in the latent space and explored the disentanglement between various semantics to achieve precise control over facial attributes.
@@
Yao et alyao2021latent ICCV’21 G.+T. By integrating explicit decoupling terms and identity-consistent terms into the loss function, the preservation of facial identity information is improved, resulting in high-quality face editing in videos.
@@
HifaFace gao2021high CVPR’21 G. Proposed a solution based on wavelet transform to address the issue of partial loss of attribute information when generating edited results due to "cyclic consistency" problems.
@@
Preechakul et alpreechakul2022diffusion CVPR’22 D. When encoding images, it is divided into semantically meaningful parts and parts that represent the details of the image.
@@
FENeRF sun2022fenerf CVPR’22 G.+NeRF ➊➍ The introduction of semantic masks into the conditional radiance field enables finer image textures.
GuidedStyle hou2022guidedstyle NN’22 G. Generating faces after face editing is guided based on facial attribute classification. The introduction of a sparse attention mecha -nism enhances the manipulation of individual attribute styles.
@@
FDNeRF zhang2022fdnerf SIG.’22 G.+NeRF The introduction of the Conditional Feature Warping (CFW) module addresses the issue of temporal inconsistency caused by dynamic information in the process of face editing in videos.
@@
AnyFace sun2022anyface CVPR’22 G. ➏➑ Proposed a dual-branch framework for text-driven facial editing, with coordination achieved between the two branches through a Cross-Modal Distillation (CMD) module.
TransEditor xu2022transeditor CVPR’22 G. ➌➊ Emphasizing dual-space GAN interaction’s importance, a transformer architecture is introduced for improved interaction.
@@
Huang et alhuang2023collaborative CVPR’23 D. Proposed the concept of assisted diffusion, integrating individual multimodalities to explore the complementarity between differe -nt modalities.
@@
Ozkan et alozkan2023conceptual ICCV’23 G. The entangled attribute space is decomposed into conceptual and hierarchical latent spaces, and transformer network encoders are employed to modify information in the latent space.
@@
CIPS-3D++ zhou2023cips TPAMI’23 G.+NeRF ➊➒ Replaced the convolutional architecture with an MLP (Multi-Layer Perceptron) architecture to achieve faster rendering speeds.
ClipFace aneja2023clipface SIG.’23 G.+3DMM Learned texture generation from large-scale datasets, enhancing generator performance through generative adversarial training.
@@
TG-3DFace yu2023towards ICCV’23 G. ➌➏ For different scenarios, two sets of text-to-face cross-modal alignment methods were designed with specific focuses.
@@
VecGAN++ dalva2022vecgan TPAMI’23 G. Orthogonal constraint and disentanglement loss are used to decouple attribute vectors in the latent space.
@@
DiffusionRig ding2023diffusionrig CVPR’23 D. 3DMM and diffusion model integration propose a two-stage method for learning personalized facial details.
@@
Kim et alkim2023diffusion CVPR’23 D. Proposed a method for facial editing in videos based on the diffusion model.
@@
SDGAN et alHuang2024SDGAN AAAI’24 G. SDGAN introduces a semantic separation generator and a semantic mask alignment strategy, achieving satisfactory preservation of irrelevant details and precise attribute manipulation.
@@
FaceDNeRF zhang2024facednerf NIPS’24 D.+NeRF Creating and editing facial NeRFs with single-view images, text prompts, and target lighting.